1. Let’s say there are two goods X and Y (respectively draw in x-axis and y-axis). Assume, X is

a normal good and Y in an inferior good. If price of X is increased, show the income and

substitution effect in the graph and explain.

2. Evaluate whether it violated or justify WARP conditionality: A consumer makes the

following choices (when you develop the matrix, show the calculation process for the first row)

a. At prices (p1,p2)=($4,$4) the choice was (x1,x2) = (20,4).

b. At (p1,p2)=($2,$2) the choice was (x1,x2) = (10,10).

c. At (p1,p2)=($2,$4) the choice was (x1,x2) = (10,8).


3. Given the Utility Function is U = X

0.7Y

0.3 and Budget is taka 300. The original price was


(Px, Py) = (2, 2) and the new price is (Px’, Py) = (4, 2).

(a) Draw an angle curve for X by using the above information [Labelling is must] [5]

(b) Calculate the value of Compensation variation (CV) and Equivalent variation (EV)? [10]

[Hint: Use the knowledge of chapter 5 and 6 to get the solutions for ordinary demand curve]


1
Expert's answer
2021-08-13T13:57:46-0400

1:




As shown in the diagram, increase in income will shift the demand to normal good x since one can afford the expensive goods. When prices for normal goods decrease, one will shift the demand to normal good x.

2:



Calculation of the first row; At the price ($4,$4) and choice (20,4) ; 20x4 + 4x4 = 16

At the price ($4,$4) and choice (10,10) ; 4x10 + 4x10= 80

At the price ($4,$4) and choice (10,8) ; 10x4 + 4x8 = 72

The numbers in bold and diagonal are the numbers representing the cost of purchasing the bundles

while the numbers other than these shows the amount that the consumer c=would have spent if he does not purchase these are the bundles 

 

When bundle A is purchased, B and C are affordable but when B is purchased bundle A is not affordable and also, when Bundle C is purchased, Bundle A is not affordable

In situation A, Bundle A is directly revealed preferred over B and C

In situation B, Bundle B is directly revealed preferred over C

in situation C, Bundle C is directly revealed preferred over B and A 

Thus, this data does not violate WARP


The data does not violate WARP


3:

a)

Given

"U=X^{0.7}Y^{0.3}.......(1)"

Budget=300

"(P_X,P_Y)=(2,2) \space and \space (P_x^`,P_Y^`)=(4,2)"


Budget line:"P_XX+p_YY=300....(2)"

Slope of budget line "=\frac{-p_X}{p_Y}"

At price px=2,py=2

Slope of indifferent curve is MRS

"MRS=\frac{-MU_X}{MU_Y}"

Where MUx is marginal utility of X and MUy is marginal utility of y.

Slope of indifferent curve"=\frac{0.7X^{0.3}Y^{0.3}}{0.3X^{0.7}Y^{-0.7}}"

"MRS=\frac{-7Y}{3X}"

Utility optimizing condition

MRS=slope of budget line

"\frac{-7Y}{3X}=-1\\Y=\frac{3X}{7}.....(3)"

Equation 3 is the equation of the angle curve of income consumption curve.it shows the optimal consumption bundle chosen at various level of income.






b)

Put value of Y into equation 3 and 2

"P_XX+P_Y(\frac{3X}{7})=300\\X(P_X+\frac{3}{7}P_Y)=300\\X=\frac{300}{P_X+\frac{3}{7}P_Y}"

"X(P_X,P_Y)=(\frac{300}{P_X+\frac{3}{7}P_Y}).....(4)"

And

"Y(P_X,P_Y)=\frac{3}{7}×(\frac{300}{P_X+\frac{3}{7}P_Y})\\Y(P_X,P_Y)=(\frac{900}{7P_X+{3}P_Y}).....(5)"

Now

"v(P_X,P_Y)=(\frac{300×7}{7P_X+3P_Y})^{0.7}(\frac{300×3}{7P_X+3P_Y})^{0.3}"

"v(P_X,P_Y)=(\frac{300×7}{7P_X+3P_Y})×7^{0.7}×3^{0.3}"

Now compensating variation (CV)

Utility before price change

"v(P_X,P_Y)=v(2,2)=\frac{300}{20}×7^{0.7}×3^{0.3}\\v(2,2)=15×7^{0.7}×3^{0.3}.........(6)"

Indirect utility function is given by

"v(P_X,P_Y,m)=\frac{m}{7p_x+3p_y}×7^{0.7}×3^{0.3}"

Now we will find value of m for which we can get same level of utility equal to equation (6)

"\frac{M^`}{7P^`_X+3P_Y}{7^{0.7}×3^{0.3}}=15×7^{0.7}×3^{0.3}"

"\frac{M^`}{7×4+3×2}=15\\m^`=15(28+6)\\m^`=510"

So compensating variation

"CV=m^`-m\\=510-300\\=210"

Now

Equivalent variation

Utility after the price change with income equal to 300

"v(P_X,P_Y,300)=\frac{300}{28+6}(7^{0.7}×3^{0.3})"

To get utility equal to "v(P_X,P_Y,300)"

Income would have to be

"\frac{300}{34}(7^{0.7}×3^{0.3})=\frac{m^{"}}{7P_X+3P_Y}(7^{0.7}\u00d73^{0.3})\\\\\\frac{300}{34}=\\frac{m^{"}}{7×2+3×2}\\m^{"}=300\u00d7\\frac{20}{34}\\\\m^{"}=176.47"

Now equivalent variation

"EV=300-m^{"}\\=300-176.47\\=123.53"